OpenStudy (anonymous):

Solve 2x+18y=-9 4x+18y=-27

OpenStudy (jhannybean):

you have two equations regarding y, solve one equation for y, and substitute it into the other

OpenStudy (anonymous):

subtracting equation 1 from equation 2 works too

OpenStudy (jhannybean):

\[2x+18y=-9\]\[18y=-2x-9\]\[y=-\frac{1}{9}x-\frac{1}{2}\] use this equation in \(4x+18y=-27\)\[4x+18(-\frac{1}{9}x-\frac{1}{2})=-27\]\[4x-2x-9=-27\]\[2x-9=-27\]\[2x=-18\]\[x=-9\] plug x back into any equation to solve for y. \[2(-9)+18y=-9\]\[-18+18y=-9\]\[18y=9\]\[y=\frac{1}{2}\]